Today, I added a new out-of-loop transmission PD (Thorlabs PDA10CS) for the north path. This will be helpful in future measurements of RIN coupling to beatnote noise. This PD is added at (8, 42) using the dumped light. The optical layout would be updated in a few days. I've also connected Acromag channel for North Transmission DC to this photodiode, so the transmitted power channels and the mode matching percentage channel of North Cavity are meaningful again.
ISS Gain for the Northside has been increased to 2x10000 since half of the light is now being used by the OOL PD.
Today, I added a new out-of-loop transmission PD (Thorlabs PDA10CS) for the south path. This will be helpful in future measurements of RIN coupling to beatnote noise. This PD is added at (1, 40) using the dumped light. The optical layout would be updated in a few days. I've confirmed that this photodiode is reading the same RIN as read earlier in CTN:2555. I've also connected Acromag channel for South Transmission DC to this photodiode, so the transmitted power channels and the mode matching percentage channel of South Cavity are meaningful again.
We realized that the half-wave plates before the EOAMs probably had no real function in the setup and therefore we proceeded to remove the one from the north path at (39,121) aka row 39, column 121 of the Optical layout.
After this was done, we had to re-adjust the quarter wave-plate (39,112) after the EOAM (39,115) to make sure that the EOAM was still functioning about the 50% transmission point. The beam going into the PMC was also re-aligned by adjusting the two mirrors at (32,92) and (37,92). Finally, the mirror at (43,88) was adjusted to align the beam reflecting from the PMC into the photo-diode.
We were able to re-lock the north PMC and north cavity after increasing the power in that path by adjusting some waveplates.
As may be expected, the sign of the ISS feedback had to be inverted. The ISS actuates on the EOAM; removing the half-wave plate would have switched the circularity of the polarization of the beam entering the PBS at (39,110), so the sign of the voltage that would have previously caused the transmission to increase would now cause it to decrease and vice versa.
I think you have to protect this SimPlant board from the HV from the FSS board ?
I added a capacitor to c8 in the FSS plant D2000020. This yielded the transfer function shown below. This seems to have pushed the pole lower.
After testing with Anchal on the FSS box we found that this seems to have done the trick. More to follow.
We added hooked up the plant and measured the in loop transfer function (shown below). The circuit locked in the system... so it works.
Next, we want to measure it with the south FSS to compare the two since anchal tells me he has never seen a direct comparison of the two.
We also need to calculate the gain of the full system. I am trying to do this on my own and get help from Anchal when I run into a problem.
# SR785 Measurement - Timestamp: Mar 10 2020 - 17:51:29
# Parameter File: TFSR785template.yml
#---------- Measurement Setup ------------
# Start frequency (Hz) = 100000.000000
# Stop frequency (Hz) = 10.000000
# Number of frequency points = 300
# Excitation amplitude (mV) = 15.000000
# Settling cycles = 1
# Integration cycles = 25
#---------- Measurement Parameters ----------
Last night I switched off all the fans in the lab and we have reached the lowest ever recorded beatnote noise.
Latest BN Spectrum: CTN_Latest_BN_Spec.pdf
Daily BN Spectrum: CTN_Daily_BN_Spec.pdf
CTN: 2551 : Comparison between Out of loop vs In loop RIN
I took spectrum of Out-of-loop (OOL) photodiode and In-loop (IL) photodiodes with transmitted light from the cavities when ISS is on in both paths at gain value of 2x10000.
These measurements take a long time as I take a median over 500 single measurement instances. I was able to increase gain later after having taken in-loop noise measurement, so I didn't repeat it. But what I can do is a take a quick in-loop measurement today with simple averaging and no error bars and post it here for comparison. If we are really interested, I can run overnight measurement for in-loop at this gain as well.
weird - why is the Gain different for in loop and out of loop ?
Attached are the latest transmitted RIN measurements.
Anchal and I connected the circuit to the EOM and PZT in a closed loop (LIGO-D040105). We looked at the output as seen in the Oscilloscope data and it does not look like the sinusoid that I was expecting. It looks like the sum of a square wave and a misshapen sinusoid. I think that there is some sort of reflection of the signal that is causing this but I don't know where. When we took the spectrum of that data and we saw a large peak at around 2kHz and then harmonics of the signal. (Note there are two spectrums both with the same data just one is on a log-log scale)
We also measured the open-loop transfer function of the circuit and the results are given below. The unity gain frequency is 1240 Hz. This is different from the 2688 Hz that is the location of the first peak in the spectrum. We had thought that maybe these harmonics were a result of the UGF but they don't occur at the same frequency.
We will try to figure out why we have harmonics and what we can do to prevent them. We really have no idea what they are.
Measurement at 3 am in the morning today has been the lowest ever recorded beatnote noise. The lasers have been locked for more than a week and the temperature of the cavities is also very. The ISS gains were increased yesterday to 5x1000 on each loop. I've also added RIN measurement and implied photothermal noise.
The issue isn't visible right now. I'm not sure what changed but with a similar power level, I'm able to increase the gain on North ISS to much higher than before. Currently, I'm taking RIN measurements (still using in-loop detector) to update the noise budget plot which is taking some time as I'm trying to measure the uncertainty in the measurement as well. So, the step response plot would come later.
The beatnote detector SN101's DC output has electronically railed. This might be due to a disconnection inside which may or may not be intentional. I don't think I should meddle around with that detector so close to the result. I will instead replace the other 1611 detector in the blocked port of beam splitter (11,19) with a thorlabs PDA10CS and use that for RIN measurement. But since, the loops were in modd of being stable today, I decided to go ahead and take the measurement with current settings. This will also work as in-loop measurement for comparison later.
do some step response and swept sine and post plots
I'm currently struggling with a problem in North ISS which I think needs to be documented here. Here's the synopsis:
I would like it if anyone has any comments on any of the points above or suggestions to tackle this problem. I can make some measurements and post them on requests.
After installing the new ISS, the noise is even lower in lower frequencies. Still no change in the noisy peaks at 400 Hz and around 800 Hz. There is also no difference in noise floor above 200 Hz.
Relevant elog post:
CTN: 2542: Installed New ISS on both paths using SR560s
I've installed new ISS consisting of one SR560 per path.
After switching out a bad opamp we have clearer results.
Next steps: moving the real value closer to the predicted value and determine the voltage the circuit can handle.
After installing the preliminary ISS, which I'll change tomorrow as per Rana's suggestions, we see some reduction in the beatnote noise in the lower frequency region. I think I should also have an estimate curve for the coupling of laser intensity noise into the final result. I can maybe make some sort of transfer function measurement from actuation on intensity to the beatnote frequency itself using moku.
CTN: 2538 : Installed ISS on both paths using SR560s
CTN:2539 : Rana's suggestion on ISS.
you don't need 6(!!) SR560s - just 1 for each loop:
Then use multi-res spectra and check out the out-of-loop noise (with loop on/off). The in-the-loop noise is always an underestimate.
I've installed ISS on both paths using 3 SR560s each. Preliminary feedback is setup to get a stable loop. More optimization with TF analysis can be done further.
The PCB came in and I have assembled it. The preliminary look at the transfer function on the AG4395A shows that there is a problem. The function is very noisy and the signal is very low. I will go through and verify all the solder points are fully connected and generally debug the circuit. A quick measurement of the transfer function on the SR785 of the low frequency (10Hz-100kHz) showed good results earlier so I fear something has come loose. Note the low-fq measurement was not recorded properly. so I have no graph for it.
I think this is a small fix that needs to be made.
The graphs below show the EOM and PZT path TF measurements on the AG4395A. They make it clear that something is not connected. This is just a progress report.
With some of the changes done recently, the beatnote noise has lowered in the low-frequency region indicating reduction in scatter.
CTN:2530 : Increased sampling rate to 125kSa/s; lowest noise in higher frequencies
CTN:2533 : blocked NF1811 with hex beam dump
CTN:2535 : BN Detector was saturated. Reduced laser powers.
CTN:2531 : Further iterated back and forth to optimized FSS Gains.
I measured dark noise of the beatnote detector reaching moku and its effect on measured beatnote frequency noise.
I have blocked the unused output port of the beam splitter before our beatnote detector with a hex beam dump at (13, 19). This was being used for broadband detector NF1811. We don't need it now.
I found that the beatnote detector was actually saturating and the output was not a good pure sinewave. I've reduced the laser powers reaching the intensity to avoid that so that 20 dB coupled output of beatnote remains around 200 mVpkpk. Following is the summary of changed settings:
However, the beatnote did not change because of these changes showing that moku is strictly sensitive to zero crossings of the acquired signal rather than its shape near the edges.
I took beatnote measurements and spanned gain values in the PMC to see if stability issues in PMC loops can be affecting the FSS downstream.
I have increased the sampling rate of moku to 125 kSa/s (fastest allowed) for the frequency-time series acquisition. After reading matermost chat of Sean Leavey etc, I felt I might have downconverted aliased noise as earlier sampling rate was just 15.625 kSa/s. This didn't change the noise below 1 kHz but we see some improvement above 1 kHz so I'll keep this from now on. The measured time series files are not around 900 Mb, so I'm not saving them and only keeping the psd calculated using modifiedPSD.py script.
CTN:2521 and replies: Beanote Spectrum vs FSS Gain Values
CTN:2528: Cleaned up table; Installed hex beam dumps
The PMC error signals have some weird broadened oscillations in them:
Above, pink is North and green is South PMC Error Signal taken from Mixer Out port.
Today I cleaned up the table, removed Scott's RFAM measurement setup and installed hex beam dumps on the input rejection of faraday isolators.
Board Plant Circuit should come in on Feb. 18th. It is really taking its time.
I took beatnote measurements and spanned gain values (COM and FAST) on South side to see the variation in beatnote with them.
Yesterday I took beatnote measurements and spanned gain values (COM and FAST) to see the variation in beatnote with them.
Today we have measured the lowest beatnote spectrum till now. This happened because I set the FSS gain values to lower than the maximum I could reach.
Relevant Elog Post:
CTN:2522: Beanote Spectrum vs FSS Gain Values
CTN:2518: NFSS Boost Stage
CTN:2514: SN010 (South RFPD) Notch Improved
CTN:2512: SN009 (North RFPD) Notch Improved!
In a discussion with Craig sometime back, it was brought up what happens when I lower the gains of the FSS loops. So today I did a test which lowers the Common and Fast Gain values on the FSS boxes by 3 dB in each step and sees what happens to the beatnote.
Added all documents to the LIGO document for the FSS Plant Circuit LIGO-D2000020-v1. The updated board and schematic are also below with the zip file of all the Eagle files.
The board was ordered from JLCPCB. The current status is "In Production." Its production will take longer because of the Chinese new year but its shipping should not because it is being shipped with DHL. The low-noise thin-film resistors were ordered from Digikey and should arrive on Friday 01/31/2020.
The next step is assembling the circuit and putting it in a box. I can't assemble the circuit until the parts arrive but I can get a box and add the power supply. This will make the production faster once the board arrives.
Once the board is assembled its transfer functions will be measured and I will work with Anchal to implement it. We are almost done.
After the Notch improvement, I took 20 dB coupled outputs of the RF out ports of the FSS RFPDs, SN009 (North) and SN010 (South), when cavities were locked or unlocked.
CTN:2470 FSS Diagnostics - RFPD RF Ouput under inspection
Today Koji came to the lab to help me out with the FSS and give me few tips.
We made some changes to the boost stage at U7 in North FSS Box board D040105-C. All the changes were made by Koji personally. He replaced Koji showed me how to solder wires and SMD components so that I can do it better next time. We soldered the wires for the boost switch, replaced the resistor R29 with a thin film 5.6 kOhm resistor and replaced the capacitor (however the older capacitor was fine too). Overall, we made the arrangement of the components neater in the stage.
We figured that the only way to measure the transfer function of this stage when boost is on is to provide some reasonable offset as well so that the opamp doesn't saturate at DC. Hence, I do not have a good measurement of the stage before the changes were made. But after the touch-up, the stage is very close to the expectation as shown in the plot.
Relevant elog posts:
CTN:2384 : Mentioned that something is wrong with the boost stage. But It was just the issue of measuring it wrong.
Repeated these measurements with maximum possible gain values in the FSS loops.
Parallel relevant threads:
CTN:2514 : FSS Diagnostics - SN010 (South RFPD) Notch Improved
CTN:2516 : Notch improved on FSS RFPDs. Gain values increased as well.
South Common Gain: 24 dB , Fast Gain: 18 dB
North Common Gain: 14 dB, Fast Gain: 14 dB
I incrased the above values as much as I could without getting oscillations in the loop.
CTN:2517: TTFSS OLTFs with Maximum Gain
Repeated these measurements after changes to the RFPDs (Notch improved).
I made the following changes to SN009:
CTN:2517: TTFSS OLTFs with Maximum Gain
I also designed a board for the circuit. This is based on the Zero model from the previous post.
(These designs have not been given a LIGO DCC number yet so I haven't added one to the design)